# Cosmological effect of coherent oscillation of ultralight scalar fields in a multicomponent universe

### v1

@@ IV Verification and applicability of our method in the spatially flat expanding FLRW spacetime
(d) Evolution of averaged out dark matter equation of state (EoS) parameter ($\omega$) with redshift (z) for different dark energy models.
@@
(f) Evolution of the effective equation of state (EoS) parameter ($\omega_{\text{eff}}$) of the system with redshift (z) for different dark energy models.
@@
Figure 1: Evolution of cosmological variables as a function of redshift in a universe composed of radiation (R), baryonic matter (B), dark matter ($\phi$), and dark energy. The horizontal axis shows $\text{Log}_{10}(1+\text{z})$, where z is the redshift, plotted in decreasing order to represent the forward progression of cosmic time. The dark energy component is modeled using three distinct approaches: (1) a cosmological constant with $\Lambda=1.08\,u^{2}$; (2) a quintessence-like scalar field ($\psi$) with an exponential potential $V(\psi)=V_{0}e^{-\lambda\psi}$, with $V_{0}=1.38\,u^{2}$; and (3) a phantom-like scalar field ($\psi$) with the same potential form but $V_{0}=0.91\,u^{2}$. In both scalar field models, the parameter $\lambda=1$. The dark matter component is represented by an oscillating scalar field $\phi$ with mass $m=1.93\times 10^{11}\,u$. All quantities are expressed in geometrized units with $u=10^{-28}\leavevmode\nobreak\ \mathrm{cm}^{-1}$.
We now analyze the evolution of a multi-component universe consisting of radiation (R), matter (M), and dark energy (DE), beginning from an early radiation-dominated epoch and continuing up to the present day, characterized by a scale factor $a_{0}=1$. The matter content includes both baryonic matter (B) and cold dark matter (CDM), where cold dark matter is modeled by the oscillating scalar field $\phi$ while the baryonic matter is treated as pressureless dust with its energy density scaling as $1/a^{3}$. To ensure radiation domination at early times, we set initial conditions at a scale factor $a_{i}=2.24\times 10^{-4}$, corresponding to a time well before the epoch of radiation–matter equality. Here, the scale factor will be dimensionless because the universe is flat. At this initial time, $t=t_{i}$ the radiation energy density is $\rho_{(\text{R})_{i}}=\rho_{(\text{R})_{0}}/a_{i}^{4}=5.56\times 10^{10}\leavevmode\nobreak\ u^{2}$, the baryonic matter density is $\rho_{(\text{B})_{i}}=\rho_{(\text{B})_{0}}/a_{i}^{3}=7.02\times 10^{9}\leavevmode\nobreak\ u^{2}$ , and the dark matter energy density is $\rho_{(\text{CDM})_{i}}=\rho_{(\text{CDM})_{0}}/a_{i}^{3}=3.74\times 10^{10}\leavevmode\nobreak\ u^{2}$. Since dark matter is modeled using the oscillating scalar field $\phi$, we determine its initial conditions to ensure that its averaged out energy density $\langle\rho\rangle$ given by Eq. (34) matches $\rho_{(\text{CDM})_{i}}$ at $t=t_{i}$. Specifically, we choose: $\phi_{+}(t_{i})=\phi_{-}(t_{i})=10^{-6},$ $\dot{\phi}_{+}(t_{i})=\dot{\phi}_{-}(t_{i})=-\frac{3H_{i}}{2}\times 10^{-6}\leavevmode\nobreak\ u$, in accordance with Eq. (50). Here, $H_{i}=\left.\frac{\dot{a}}{a}\right|_{t=t_{i}}=\frac{\dot{a_{i}}}{a_{i}}=1.82\times 10^{5}\leavevmode\nobreak\ u$ denotes the Hubble parameter at the initial time. Despite the large initial value of the Hubble parameter, the condition $m\gg H$ remains satisfied throughout the evolution considered here, thus validating the applicability of our method.
For dark energy, we have considered two distinct models:
  1. 1.
    Cosmological Constant ($\Lambda$): Representing a constant vacuum energy density $\rho_{(\Lambda)}$.
  2. 2.
$\left(\frac{\dot{a}}{a}\right)^{2}=\frac{1}{3}\left[\langle\rho\rangle+\frac{\rho_{(R)_{0}}}{a^{4}}+\frac{\rho_{(B)_{0}}}{a^{3}}+\rho_{(\Lambda)}\right]\,,$ (55)
The resulting cosmological evolution is shown in Fig. [1]. Numerical analysis confirms that the coherently oscillating ultralight scalar field closely replicates the dynamical behavior of the CDM model, with the averaged out scalar field EoS parameter, $\omega$, remaining effectively zero throughout. In Fig. [1(a)] we plot the energy density parameters of the individual components as, baryonic matter (B), radiation (R), scalar field dark matter ($\phi$) and the cosmological constant ($\Lambda$) as a function of ${\rm Log}_{10}(1+\text{z})$, where the redshift z, plotted in decreasing order to represent the forward progression of cosmic time. This visualization effectively captures the time evolution of the density parameters for each component.
The evolution of the multicomponent universe follows the standard cosmological sequence. It begins in a radiation-dominated phase. As expansion proceeds, the matter energy density surpasses the radiation density, marking the onset of the matter-dominated era. Eventually, dark energy, represented by the cosmological constant, dominates and drives late-time accelerated expansion. Importantly, the oscillating ultralight scalar field used to represent dark matter replicates the standard behavior. It reproduces the correct redshift values for both the radiation–matter equality and the matter–dark energy equality, consistent with what is expected in the case of cold dark matter. The effective EoS in Fig. [1(f)] evolves accordingly, starting from a positive value during the radiation era and gradually approaching $-1$, as the system transitions into the dark energy-dominated phase.
The Fig. [1] illustrates the results when the dark energy is modeled by quintessence and phantom fields. The numerical results in Figs. [1(b), 1(c), 1(d)] show a broadly similar behavior to the case of the cosmological constant case as expected; some subtle differences appear in the evolution of the effective equation of state (EoS) parameter, $\omega_{\text{eff}}$, in Fig. [1(f)], which now deviates from the value it had in the cosmological constant model and evolves differently depending on the chosen scalar field model. Also, the dark energy EoS parameter in Fig. [1(e)] is no longer strictly constant, as it is for a cosmological constant, but evolves with time, taking values greater than $-1$ for quintessence and less than $-1$ for phantom fields.
IV.2 The case when CDM, radiation, and dark energy are present and the CDM sector has non-minimal coupling with the radiation sector
In the non-minimally coupled case, where the scalar field decays into radiation, we begin with the same initial values as in the previous subsection, corresponding to a radiation domination epoch. A decay constant of $\Gamma=10^{-4}\leavevmode\nobreak\ u$ is introduced to govern energy transfer from the scalar field to radiation. In this scenario, the dark matter component, represented by the oscillating ultralight scalar field, gradually decays into radiation in the presence of baryonic matter and dark energy. We first analyze the evolution of a system where dark energy is modeled by the cosmological constant, and then proceed to the case where dark energy is described by a dynamical scalar field.
(d) Evolution of dark matter equation of state (EoS) parameter ($\omega$) with redshift (z) for different dark energy models.
(f) Evolution of the effective equation of state (EoS) parameter ($\omega_{\text{eff}}$) of the system with redshift (z) for different dark energy models.
@@ IV.1 The case when the CDM sector, baryonic matter, radiation, and dark energy are present
Figure 2: Evolution of cosmological variables as a function of redshift in a universe composed of radiation (R), baryonic matter (B), dark matter ($\phi$), and dark energy, where radiation is non-minimally coupled to dark matter with coupling constant $\Gamma=10^{-4}\leavevmode\nobreak\ u$. The horizontal axis shows $\text{Log}_{10}(1+\text{z})$, where z is the redshift, plotted in decreasing order to represent the forward progression of cosmic time. The dark energy component is modeled using three distinct approaches: (1) a cosmological constant with $\Lambda=1.08\,u^{2}$; (2) a quintessence-like scalar field ($\psi$) with an exponential potential $V(\psi)=V_{0}e^{-\lambda\psi}$, with $V_{0}=1.38\,u^{2}$; and (3) a phantom-like scalar field ($\psi$) with the same potential form but $V_{0}=0.91\,u^{2}$. In both scalar field models, the parameter $\lambda=1$. The dark matter component is represented by an oscillating scalar field $\phi$ with mass $m=1.93\times 10^{11}\,u$. All quantities are expressed in geometrized units with $u=10^{-28}\leavevmode\nobreak\ \mathrm{cm}^{-1}$.
The time evolution of this non-minimally coupled system for the first case is governed by Eqs. (44), (45), (46), along with the modified Friedmann equation:
$\left(\frac{\dot{a}}{a}\right)^{2}$ $=$ $\frac{1}{3}\left[\langle\rho\rangle+\rho_{(R)}+\frac{\rho_{(B)_{0}}}{a^{3}}+\rho_{(\Lambda)}\right]\,.$ (57)
The resulting time evolution of the various cosmological variables from this coupled system of equations is shown in Fig. [2], where the oscillating ultralight scalar field dark matter decays into radiation under the influence of a cosmological constant. Despite the decay, the effective EoS parameter of the dark matter component remains that of pressureless dust. Importantly, the density parameters of the individual components evolve such that their present-day values are consistent with observations, indicating that even a small coupling between dark matter and radiation can yield a viable cosmological evolution. Our results show that standard cosmological dynamics can admit a small value of $\Gamma\sim 10^{-37}$ eV. This small decay rate from the CDM particles to photons does not affect the results of cosmological dynamics till now. We will come back to this topic at the end of this subsection. Next, we present the results assuming a dynamic dark energy sector.
$\left(\frac{\dot{a}}{a}\right)^{2}$ $=$ $\frac{1}{3}\left[\langle\rho\rangle+\rho_{(R)}+\frac{\rho_{(B)_{0}}}{a^{3}}+\rho_{(\psi)}\right]\,.$ (58)
Fig. [2] illustrates the time evolution of the various cosmological variables for the system described. The behavior exhibited closely parallels that observed in the scenario where the cosmological constant models dark energy. As in the minimally coupled case, both the dark energy equation of state (EoS) driven by the scalar field and the overall effective EoS of the system depend on the type of scalar field used to model the dynamic dark energy sector. The results shown in the above plots are satisfactory, showing that the generalization of the coherent oscillation model of the ultralight scalar field model works properly in the cosmological sector. In this paper, we have not compared our results with modern cosmological data; in a future publication, we would like to do that and set a more accurate phenomenological bound on $\Lambda$. In the present calculations, we have used a value of $\Gamma$ which is sufficiently smaller in value when compared to the Hubble parameter value. Technically, we are in the weak dissipative regime where the scalar field energy weakly transforms into radiation. In the far future, the value of the Hubble parameter will reduce and then $\Gamma\sim H$, and then the above results may change. If such a regime at all arises, then we can have conversion of the whole CDM energy into radiation.
(f) Evolution of the effective equation of state parameter ($\omega_{\text{eff}}$) of the system with normalized time ($t/t_{\text{v}}$) for different dark energy models.
(g) Evolution of matter equation of state parameter ($\omega$) with normalized time ($t/t_{\text{v}}$) for different dark energy models.
Figure 3: Time evolution of system variables for an overdense region treated as a closed system with positive curvature ($k=1$). The radiation (R) is non-minimally coupled with a coupling constant $\Gamma=10^{-4}\leavevmode\nobreak\ u$ to an oscillating ultralight scalar field ($\phi$) representing dark matter, with mean energy density $\langle\rho\rangle$. Dark energy is incorporated through three distinct models: (1) a cosmological constant with $\Lambda=1.08\leavevmode\nobreak\ u^{2}$; (2) a quintessence-like scalar field ($\psi$) with potential $V(\psi)=V_{0}e^{-\lambda\psi}$, where $V_{0}=1.38\leavevmode\nobreak\ u^{2}$; and (3) a phantom-like scalar field ($\psi$) with the same potential form but $V_{0}=0.91\leavevmode\nobreak\ u^{2}$. For both scalar field models, the parameter $\lambda=1$. All variables are evaluated up to the virialization time $t_{\text{v}}$. All quantities are expressed in geometrized units with $u=10^{-28}\leavevmode\nobreak\ \mathrm{cm}^{-1}$.
V Application of our method in gravitational collapse
In this section, we show that the generalization of the coherent oscillation of the ultralight scalar field model can behave like the CDM sector even when we are dealing with gravitational collapse. Such a collapsing scenario was recently studied in Ref. Saha:2024irh where the authors used an approximated version of the coherent oscillation of the ultralight scalar field model. Unlike the previous work, in the present article, we will use the exact generalization of the model with the proper initial conditions.
Here we investigate the dynamics of gravitational collapse in a system where dark matter is non-minimally coupled to radiation in the presence of dark energy. The evolution is studied in a closed FLRW spacetime with positive spatial curvature $(k=1)$. Dark matter is modeled by the coherently oscillating ultralight scalar field. We focus on the behavior of a small, spherically symmetric overdense region embedded in the expanding universe. Due to its initially higher density relative to the background, this region gradually decouples from the Hubble flow, experiences a slowdown in expansion, and eventually reaches a turnaround point, initiating its gravitational collapse. Our analysis, however, does not proceed to the final singularity of the collapse. Instead, it concentrates on the stage when the system attains virial equilibrium and stabilizes. The primary motivation for introducing a phenomenological mechanism to halt the collapse is rooted in the observational evidence of large-scale structures in the universe, which are believed to form via such virialization processes. A collapsing region must, therefore, reach a virialized state at some stage in its evolution.
$\frac{R_{\text{v}}}{R_{\text{max}}}=\frac{1}{2},$
In the present case, we assume that the overdense region initially does not contain baryonic matter or radiation; radiation is generated dynamically through the decay of dark matter. Thus, we set the initial radiation density at $t=0$ to zero, $\rho_{R}(0)=0$, in principle. For numerical stability, we begin with a small seed value, $\rho_{R}(0)=0.00001\leavevmode\nobreak\ u^{2}$. As before, dark energy is modeled in two distinct ways: as a cosmological constant and as a scalar field $\psi$, in both cases, we use the initial conditions and parameter values as given in subsection III.2. This general framework allows us to study any epoch by choosing different values of the scale factor. We select $a=0.1\leavevmode\nobreak\ u^{-1}$, here $a$ is not dimensionless like before, it has dimensions of length because we are working with $k=1$. The initial conditions for the oscillating scalar field at, $t=0$ are given as: $\phi_{+}(0)=\phi_{-}(0)=10^{-10},$ $\dot{\phi}_{+}(0)=\dot{\phi}_{-}(0)=-\frac{3H_{i}}{2}\times 10^{-10}\leavevmode\nobreak\ u$, with $H_{i}=0.454\leavevmode\nobreak\ u$. All quantities are rendered dimensionless by normalizing with respect to their values at virialization: $a/a_{v}$, $t/t_{v}$, and $\rho/\rho_{v}$. The mass of the ultralight scalar field ($m$) and the coupling constant ($\Gamma$) have the same value they had in the previous sections.
$\left(\frac{\dot{a}}{a}\right)^{2}$ $=$ $\frac{1}{3}\left[\langle\rho\rangle+\rho_{\text{(R})}+\rho_{(\Lambda)}\right]-\frac{1}{a^{2}}\,.$ (59)
$\left(\frac{\dot{a}}{a}\right)^{2}$ $=$ $\frac{1}{3}\left[\langle\rho\rangle+\rho_{\text{(R})}+\rho_{(\psi)}\right]-\frac{1}{a^{2}}\,.$ (60)
VI Results and discussions
We have seen that the equivalence of the coherent oscillating ultralight scalar field model and the CDM sector allows the ultralight scalar field to have a minute decay rate $\Gamma\sim 10^{-37}$ eV. Although this decay rate does not produce a significant difference from the standard cosmological development, this small value of $\Gamma$ can have a nontrivial effect in gravitational collapse. The dark matter collapse locally will produce faint radiation. Moreover, in the future $H\sim\Gamma$ (signaling the onset of a strong dissipative regime) and consequently all the CDM will start to convert into radiation. If we allow the small $\Gamma$ to exist, then we can predict that the universe will convert all its CDM component into radiation in the future.

### v2

@@ C.1 Evolution of Cosmological Parameters when CDM decays into radiation
(d) Evolution of time-averaged dark matter equation of state (EoS) parameter ($\omega$) with redshift (z) for different dark energy models.
@@
[…]
@@
Figure 4: Evolution of cosmological variables as a function of redshift (z) in a universe composed of radiation (R), baryons (B), dark matter ($\phi$), and dark energy, where dark matter decays into radiation with decay constant $\Gamma=10^{-4}\penalty 10000\ u$. The horizontal axis shows $\log_{10}(1+z)$, plotted in decreasing order to represent the forward progression of cosmic time. Dark energy is modeled in three ways: (1) a cosmological constant with $\Lambda=1.08\,u^{2}$; (2) a quintessence scalar field ($\psi$) with exponential potential $V(\psi)=V_{m}e^{-\lambda\psi}$, $V_{m}=1.38\,u^{2}$, $\lambda=1$, and initial conditions $\psi(t_{i})=0.0001$, $\dot{\psi}(t_{i})=0.0001\penalty 10000\ u$; (3) a phantom scalar field with the same potential form and same initial and parameter values except $V_{m}=0.91\,u^{2}$. Dark matter is an oscillating scalar field $\phi$ with mass $m=1.93\times 10^{11}\,u$ and initial conditions $\phi_{\pm}(t_{i})=10^{-6}$, $\dot{\phi}_{\pm}(t_{i})=-\frac{3H_{i}}{2}\times 10^{-6}\penalty 10000\ u$, with $a_{i}=2.24\times 10^{-4}$ and $H_{i}=1.82\times 10^{5}\penalty 10000\ u$. All other cosmological parameters follow Planck 2018 Planck:2018vyg , and quantities are expressed in geometrized units $u=10^{-28}\penalty 10000\ \mathrm{cm}^{-1}$.
Here, we present the evolution of the density parameters of the individual components—radiation, baryonic matter, dark matter, and dark energy—as well as the equation-of-state (EoS) parameters for dark matter, dark energy, and the total cosmic fluid, for the case where dark matter decays into radiation with a small decay constant $\Gamma=10^{-4}\,u$, expressed in natural units as $10^{-37}\,\mathrm{eV}$. The plotted quantities in Fig. (4) demonstrate that their evolution throughout cosmic history remains essentially identical to the case without dark matter decay, with all variables following the same overall trends. These results confirm that a small coupling between dark matter and radiation, corresponding to the chosen decay constant, is cosmologically viable and does not significantly alter the background evolution of the universe. The present-day values of the density parameters remain consistent with observational constraints, indicating that even such a small decay rate is compatible with standard cosmology.
C.2 Numerical Energy-Balance Diagnostic
To test the numerical accuracy of the decay runs, we have computed the residual of the radiation continuity equation, $\mathcal{R}=\dot{\rho}_{R}+4H\rho_{R}-\Gamma\langle\dot{\phi}^{2}\rangle,$ along with its normalized form expressed as a percentage relative to $H\rho_{R}$, where the percentage uses the absolute value of $\mathcal{R}$ to quantify the maximum fractional deviation and $\rho_{R}$ is the total radiation energy density. The resulting quantities, shown in Fig. (5) for the cosmological constant, quintessence, and phantom field cases, remain consistent with numerical precision throughout the evolution. The residual $\mathcal{R}$ remains close to zero throughout the evolution with respect to redshift, in agreement with the theoretical expectation. The maximum relative residual is about $0.08\%$ for the cosmological constant model, $0.10\%$ for the quintessence like scalar field dark energy and $0.12\%$ for phantom like dark energy case, demonstrating that the total energy exchange between the scalar field and radiation is well conserved by the numerical integration scheme.
(a) Residual of the radiation continuity equation ($\mathcal{R}$) with redshift (z).
(b) Relative percentage residual of the radiation continuity equation with redshift (z)
@@ C.2 Numerical Energy-Balance Diagnostic
Figure 5: Evolution of the residual of the radiation continuity equation ($\mathcal{R}$) in $u^{3}$ units and its normalized form expressed as a percentage relative to $H\rho_{R}$, shown as functions of redshift in a universe composed of radiation (R), baryonic matter (B), dark matter ($\phi$), and dark energy. The dark matter component decays into radiation with a decay constant $\Gamma=10^{-4}\,u$. The horizontal axis shows $\log_{10}(1+z)$, plotted in decreasing order to represent the forward progression of cosmic time. The dark energy sector is modeled using three approaches: (1) a cosmological constant (black curve) with $\Lambda=1.08\,u^{2}$; (2) a quintessence-like scalar field (red curve, $\psi$) with an exponential potential $V(\psi)=V_{m}e^{-\lambda\psi}$ and $V_{m}=1.38\,u^{2}$; and (3) a phantom-like scalar field (green curve, $\psi$) with the same potential form but $V_{m}=0.91\,u^{2}$. In both scalar field models, the parameter $\lambda=1$ and initial conditions $\psi(t_{i})=0.0001$, $\dot{\psi}(t_{i})=0.0001\penalty 10000\ u$;. The dark matter field $\phi$ is modeled as an oscillating scalar field with mass $m=1.93\times 10^{11}\,u$ and initial conditions $\phi_{\pm}(t_{i})=10^{-6}$, $\dot{\phi}_{\pm}(t_{i})=-\frac{3H_{i}}{2}\times 10^{-6}\penalty 10000\ u$, with $a_{i}=2.24\times 10^{-4}$ and $H_{i}=1.82\times 10^{5}\penalty 10000\ u$. All other cosmological parameters follow Planck 2018 Planck:2018vyg , and quantities are expressed in geometrized units $u=10^{-28}\penalty 10000\ \mathrm{cm}^{-1}$.